### v1

# Surface Optimization of Aluminum Resonators for Robust Quantum Device Fabrication

Abstract
Aluminum remains the central material for superconducting qubits, and considerable effort has been devoted to optimizing its deposition and patterning for quantum devices. However, while post-processing of Nb- and Ta-based resonators has been widely explored, primarily focusing on oxide removal using buffered oxide etch (BOE), post-treatment strategies for Al resonators remain underdeveloped. This challenge becomes particularly relevant for industry-scale fabrication with multi-chip bonding, where delays between sample preparation and cooldown require surface treatments that preserve low dielectric loss during extended exposure to ambient conditions.
In this work, we investigate surface modification approaches for Al resonators subjected to a 24-hour delay prior to cryogenic measurement. Passivation using self-limiting oxygen and fluorine chemistries was evaluated utilizing different plasma processes. Remote oxygen plasma treatment reduced dielectric losses, in contrast to direct plasma, likely due to additional ashing of residual resist despite the formation of a thicker oxide layer on both Si and Al surfaces. A fluorine-based plasma process was developed that passivated the Al surface with fluorine for subsequent BOE treatment. However, increasing fluorine incorporation in the aluminum oxide correlated with higher loss, identifying fluorine as an unsuitable passivation material for Al resonators.
Finally, selective oxide removal using HF vapor and phosphoric acid was assessed for surface preparation. HF vapor selectively etched SiO2 while preserving Al2O3, whereas phosphoric acid exhibited the opposite selectivity. Sequential application of both etches yielded dielectric losses as low as $\delta_{\mathrm{LP}}=5.2\times 10^{-7}$ ($Q\mathrm{i}\approx 1.9\,\mathrm{M}$) in the single photon regime, demonstrating a promising pathway for robust Al-based resonator fabrication.
I Introduction
After decades of research, Aluminum (Al) remains a widely used material for superconducting transmons. While Nb and Ta start to replace Al in the fabrication of resonators due to their favorable properties [33, 4], Al continues to be indispensable for Josephson junction electrodes, making it the most critical component of superconducting qubits. Significant efforts have therefore been devoted to optimize Al deposition and patterning for superconducting quantum circuits [3, 11]. While post-processing strategies for Nb and Ta resonators, typically aimed to improve resonator quality with surface cleaning and oxide removal [2, 24, 10] using buffered oxide etch (BOE)[35, 1, 36], have been extensively studied, those techniques can not easily be applied to Al metallization. Systematic investigations of post-treatments for Al resonators [23, 37, 6] have proposed different strategies to address existing limitations, which are particularly important because dielectric losses associated with native and process-induced oxides on aluminum or silicon can limit coherence in superconducting circuits [12, 26].
Industrial-scale device fabrication introduces additional constraints. Multi-chip packaging, scheduling, and cleanroom throughput often prevent the immediate transfer of resist stripped samples into a cryostat. Consequently, surface treatments must remain effective even when a cooldown is delayed by several hours. Therefore, the present work focuses on surface treatments applicable to Al resonators with a controlled delay of approximately 24 h before cooldown, reflecting conditions encountered in scalable fabrication environments.
The main challenge addressed in this study is modifying the Al surface to reduce dielectric losses without damaging the Si surface in the resonator trench. Oxygen and fluorine were selected as promising passivation species. Oxygen-based treatments are compatible with the naturally forming Al2O3 layer, and resist ashing is known to reduce dielectric losses by removing polymer residues [30]. Low-energy plasma oxidation enables controlled growth and saturation of the surface oxide [19] without introducing ion-induced damage or surface roughness that could degrade resonator performance.
Fluorine-based passivation was also explored, motivated by the higher bond energy of Al–F compared to Al–O ($\mathrm{G}_{\mathrm{AlF}}=675$ kJ/mol vs. $\mathrm{G}_{\mathrm{AlO}}=502$ kJ/mol [31]), which suggests protection against re-oxidation [14, 20, 7]. The protection of the Al surface was demonstrated by the repellency of BOE, which normally attacks Al. Prior to fluorination, native oxides were removed to prepare a pristine Al surface for the fluorine plasma treatment. To date, there are no studies showing the influence of fluorine passivation on dielectric losses of superconducting Al resonators.
Finally, selective chemical post-etching was employed to remove SiO2 and Al2O3 independently. HF vapor etching selectively removed SiO2 without affecting Al2O3, whereas phosphoric acid exhibited the opposite selectivity. Since oxides are known to host two-level systems (TLS) [25, 26, 12], removing them is expected to reduce dielectric losses.
Our study demonstrated the effectiveness of surface preparation for enhancing the performance of Al-based superconducting resonators.
II Analytical Methods
X-ray photoelectron spectroscopy (XPS) measurements quantify elemental composition of surface layers. With in-situ Ar ion milling to remove the upper layers, the technique also enables depth profiling and helps to eliminate contamination such as carbon or native oxides.
In this study, we used XPS to analyze the Al, F, and O composition of CF4 plasma-treated Al chip surfaces. All samples were etched for 1 min with in-situ Ar ion milling prior to the XPS measurement to remove environmental carbon contamination.
Figure 1: Design of the resonator chips, containing nine CPW meanders with frequencies between 4 GHz and 5.6 GHz.
Cryogenic benchmarking of the post-treatments was performed at temperatures below 15 mK. The chip design consists of nine coplanar waveguides (CPW) resonators (gap: 6 µm, width: 10 µm) with evenly spaced frequencies between 4 GHz and 5.6 GHz, all capacitively coupled to a common feedline. The ground plane surrounding the resonators is patterned with vortex-pinning holes to mitigate quasiparticle tunneling [27].
The measurement procedure is as follows: after locating the resonance peaks with a coarse transmission sweep, the complex scattering parameter $S_{21}$ is recorded. Both is done using a vector network analyzer (VNA) [32, 13], which is capable of generating and analyzing microwave signals. The internal quality factor $Q_{\mathrm{i}}$ is then extracted by fitting the data according to Ref. [29]. Using the method described in [5], the VNA excitation power can be converted into photon number. The internal quality factor, defined as the inverse material loss $Q_{\mathrm{i}}=1/\delta$, can be expressed in terms of the dielectric loss [25]
$\delta=\delta_{\mathrm{LP}}\,\frac{1}{\left(1+\langle n\rangle/n_{\mathrm{c}}\right)^{\beta}}+\delta_{\mathrm{HP}},$ (1)
assuming a broad distribution of TLS within continuum approximation and $T\rightarrow 0$. Here, $\langle n\rangle$ is the photon number, $n_{\mathrm{c}}$ the critical photon number, $\beta$ an exponent, $\delta_{\mathrm{LP}}$ the low-power loss and $\delta_{\mathrm{HP}}$ the high-power loss. Since qubits operate at n ¡ 1, we target the single‑photon regime, where intrinsic TLS losses dominate and fit the data (see Fig. 2) with Eq. (1) and extract $\delta_{\mathrm{LP}}=1/Q_{\mathrm{LP}}$ from the low-power limit ($n\rightarrow 0$).
Figure 2: Exemplary plot of $\delta$ vs. photon number with measured data (blue) from a 4.36 GHz resonator and a fit using Eq. (1) (red).
This data processing was performed for each resonator individually, while the spread of $\delta_{\mathrm{LP}}$ was statistically analyzed using the interquartile range (IQR) with its median $\tilde{\delta}_{\mathrm{LP}}$. For the subsequent analysis, we used the low-power material loss to evaluate the suitability of our post-treatment methods for superconducting qubit applications.
III Experimental
The fabrication of resonator chips on 200 mm wafers follows the same process parameters, substrate selection, and cleaning procedures as detailed in [18] for the first metal layer. Following the post-etch resist stripping by $\mathrm{H}_{2}\mathrm{O}$ plasma ashing, which results in the formation of a thicker surface oxide compared to the native $\mathrm{SiO}_{2}$ and $\mathrm{Al}_{2}\mathrm{O}_{3}$, and EKC cleaning the wafer is coated with protective resist for dicing. The plasma ashing is needed to remove in-situ Cl contamination from the etching process, which degrade the Al surface. After dicing and resist removal using isopropanol and acetone, the fabrication process for the untreated reference samples is completed. It should be noted that we obtained all samples used in this study from a single 200 mm wafer to suppress wafer-to-wafer variability and enable a more controlled and comparable evaluation of process effects. Three different post-treatment approaches (see Fig. 3) were applied separately on these samples, which are in principle compatible with wafer-scale implementation.
Figure 3: Schematic representation of the different post-treatments applied to the resonator chips. Approach A utilized oxygen plasma under different conditions. In approach B, the Al surface was cleaned in-situ using Ar ion milling, followed by plasma processing with CF4. In approach C, HF vapor and H3PO4 acid were applied both individually and in combination.
III-A Passivation via Oxidation
This approach (see Fig. 3) aimed to compare various systems and processes of oxygen plasma. A remote oxygen plasma system with parameters akin to those reported in [19] was employed for durations of 20 sec and 120 sec, which enhance the oxide thickness (approximately 5 nm [19]) and oxygen density compared to native aluminum. Similar effects are known for silicon oxides [21]. Additionally, a direct oxygen plasma treatment using an Oxford Instruments PlasmaPro 80 system in RIE mode was conducted at high (290 W) and low (50 W) power, at 300 mT for 3 minutes. The low-power recipe represents a standard resist stripping process well established in the EMFT cleanroom, while the high-power process was tested to improve the ashing capabilities for enhanced surface cleaning.
III-B Passivation via Fluorination
Here, we focused on passivating the aluminum surface with fluorine while preserving a pristine silicon surface (see Fig. 3). The Oxford Instruments PlasmaPro 80 tool was employed in RIE mode due to its capability of executing the entire fabrication scheme in situ. Initially, the aluminum surface oxide was removed using Ar plasma, with process times varied for optimal oxide etching. Subsequently, CF4 plasma was applied at 60 W and 250 mT for 3 minutes to fluorinate the exposed aluminum surface. The stability of the passivation layer was then evaluated through subsequent buffered oxide etching (BOE) for 30 seconds or HF vapor etching of an equivalent of 150 nm SiO2 in the uEtch tool from KLA, which offered high selectivity to silicon and aluminum.
III-C Post-Etching
This approach (see Fig. 3) investigated the removal of $\mathrm{SiO}_{2}$ and $\mathrm{Al}_{2}\mathrm{O}_{3}$ from the surfaces. For silicon oxide removal, HF vapor was utilized again in the uEtch tool for its selectivity to silicon and aluminum. The process conditions were mentioned before (see Section III-B). For aluminum oxide etching, undiluted phosphoric acid ($\mathrm{H}_{3}\mathrm{PO}_{4}$), known from conventional aluminum wet etch chemistry [16], was employed due to its high selectivity for silicon/SiO2. In our methodology, we increased the concentration of the solution to maintain high selectivity for aluminum [15]. At a etch time of 1 minute, we were able to observe a saturation of the remaining aluminum oxide thickness with ellipsometry, which we attribute to the regrowth of native oxide in ambient air. Both treatments were applied to samples independently as well as in combination.
Upon completion of the post-treatment, a 24-hour coupling period begins until vacuum is reestablished during cooldown in the cryostat. During this time frame, the samples are bonded onto the PCB and mounted in the cryostat. The measurement of $\delta_{LP}$ finalizes the characterization process.
IV Results and Discussion
All chips used in this study were taken from the same wafer. The resulting metal layer thickness and linewidth variations are comparable to [18] and the samples can therefore be treated as geometrically identical. Each post-treatment described in this work was applied to two resonator chips, which were subsequently characterized in the dilution refrigerator. Some chips were lost due to handling errors, which is reflected by fewer data points in the figures. All chips were characterized using identical measurement procedures. By analyzing the $\delta_{\mathrm{LP}}$ of the resonators, we benchmark the suitability of each process for superconducting qubit applications.
IV-A Passivation via Oxidation
Figure 4: Low-power loss $\delta_{\mathrm{LP}}$ of Al resonators after different oxygen plasma treatments.
For an initial assessment of the process influence on material loss, oxygen plasma treatments were performed in remote and direct mode (see Section III-A). Remote oxygen plasma slightly reduced $\delta_{\mathrm{LP}}$ compared to untreated samples, with longer exposure times yielding marginally lower losses (see Fig. 4). Overall, $\tilde{\delta}_{\mathrm{LP}}$ for the 120 s remote plasma process showed a reduction of approximately 50% compared to untreated samples, although the IQR remain relatively large and partially overlapping. This is in contrast to others [28, 9], which reported on increased dielectric losses resulting from thicker oxides. The reduced loss may be attributed to oxygen radicals filling oxygen vacancies in the non-stoichiometric surface oxides or to additional ashing of polymer residues originating from the protective resist. The 50 W direct plasma process yielded similar results to the 20 sec remote plasma.
However, the high-power (290 W) direct plasma process increased $\tilde{\delta}_{\mathrm{LP}}$ beyond the upper quartile of the untreated reference. The elevated ion energies likely damage both the Si and Al surfaces, thereby increasing the dielectric losses. (AFM pictures pending)
IV-B Passivation via Fluorination
CF4 plasma treatments on Al surfaces were investigated as a potential passivation technique. Since CF4 plasmas can either etch or deposit fluorocarbon films depending on process parameters [22, 8], a short parameter study with varying plasma powers was carried out. Based on this, a CF4 process was selected that incorporated fluorine into the Al surface without depositing C–F residues on the Si surface (see Section III-B).
To efficiently remove the native Al2O3 layer prior to fluorination without damaging Si or Al, Ar ion milling with different durations was investigated. Figure 5 shows the XPS-derived atomic fractions of Al, F, and O for different surface preparation conditions. Increasing ion-milling duration results in higher fluorine content and reduced oxygen content, reaching saturation after approximately 5 min, with an F/O ratio of about 0.5. The incorporation of fluorine into the surface is energetically favored ($\mathrm{G}_{\mathrm{AlF}}=675$ kJ/mol vs. $\mathrm{G}_{\mathrm{AlO}}=502$ kJ/mol [31]), supporting the approach of reducing oxygen through fluorination.
Figure 5: Atomic composition versus Ar ion milling duration prior to CF4 plasma treatment. The surface was cleaned for 1 min using Ar ion milling before the XPS measurements were carried out.
These fluorination processes were subsequently applied to resonator chips and characterized at millikelvin temperatures. As shown in Fig. 6, the low-power loss of fluorinated samples increased compared to untreated samples. Longer ion-milling durations correlated with higher loss, consistent with the increased fluorine content in the Al surface layer. An increase in dielectric loss due to Ar milling is not expected, as a previous study [34] demonstrated that Ar-milling-based processes, under conditions of enhanced oxygen removal, do not result in increased internal losses in aluminum resonators.
Figure 6: Low-power loss $\delta_{\mathrm{LP}}$ of fluorinated samples for different Ar ion-milling durations.
To evaluate whether fluorine incorporation is directly responsible for the increased loss, additional post-etching using BOE or HF was performed in order to remove potential Si trench damage or amorphization [24, 17] caused by Ar ion milling [30]. For the 3 min milled sample (see Fig. 7,a), BOE did not reduce $\delta_{\mathrm{LP}}$ compared to the non-etched sample, as both IQR nearly overlap. Interestingly, BOE did not visibly damage the Al surface, which would be the case for untreated Al (see Fig. 7,b). Here, the BOE etches at the grain boundaries until the Si is reached. This suggests that the fluorinated surface is sufficiently passivated to withstand BOE, with fluorine contents as low as 27% appearing sufficient. Likewise, HF treatment did not lead to a reduction in $\tilde{\delta}_{\mathrm{LP}}$, as the IQR remains comparable.
[…]
Figure 7: (a) Effect of BOE and HF post-etching on fluorinated Al resonators. (b) If no passivation is apparent, the Al surface is visibly attacked by 30 sec BOE.
These findings suggest that fluorine incorporation stabilizes the Al surface and passivates it against aggressive Al/Al2O3 etchants. However, the low-power loss in the superconducting regime increases by a factor of approximately 2–3.
IV-C Post-Etching
As a final post-processing approach, HF and phosphoric acid were used to selectively remove SiO2 and Al2O3 from the resonator surfaces (see Section III-C) of untreated samples. The chips were subsequently cooled down for cryogenic characterization.
Compared to untreated samples, the phosphoric acid etch reduced $\tilde{\delta}_{\mathrm{LP}}$ by approximately a factor of 2 (see Fig. 8), separating both IQR. HF etching produced a similar improvement. Applying both etches sequentially further reduced the loss by a combined factor of three, resulting in $\tilde{\delta}_{\mathrm{LP}}=5.7\cdot 10^{-7}$ with $\delta_{\mathrm{LP}}=5.2\cdot 10^{-7}$ for the lower and $\delta_{\mathrm{LP}}=7.0\cdot 10^{-7}$ for the upper quartile. This corresponds for the lower quartile to a quality factor of $Q_{\mathrm{LP}}=1.9$ M, which represents the highest value observed throughout this study. These findings support the well-established idea that surface oxides host TLS defects, which increase low-power losses [25, 26, 12]. Notably, we were able to effectively reduce oxide-related losses despite a 24 h delay between fabrication and cooldown.
Figure 8: Low-power loss after HF, phosphoric acid, and combined post-etching treatments.
V Conclusion
In this study, several strategies were investigated to post-treat the surfaces of aluminium resonators prior to a cooldown occurring 24 h after processing. Oxygen and fluorine plasmas were used as passivation approaches. Plasma oxidation marginally reduced $\delta_{\mathrm{LP}}$ when low-power direct plasma or remote plasma was applied, which is likely attributed to additional ashing that removes residual resist or polymer contamination. In contrast, high-power oxygen plasma increased the losses, which can be caused by the surface impact of high energetic ions.
Fluorine plasma treatment in combination with Ar ion milling showed an increased fluorine concentration on the Al surface up to F/O $\approx$ 50% with increasing milling duration, which correlated with an increase in $\delta_{\mathrm{LP}}$. The fluorine-treated surface exhibited a passivating effect against BOE etching, enabling controlled preparation of the aluminium metallization for subsequent BOE post-treatments. This is particularly relevant for device architectures in which Nb or Ta are used in combination with Al. The fluorinated Al surfaces, including Josephson junction electrodes, can be sufficiently stabilized to allow post-BOE dipping of the entire chip, which is crucial to achieve high quality factors with Nb or Ta resonators.
Finally, selective post-etching with HF and phosphoric acid was employed to remove SiO2 and Al2O3 from the resonator surfaces. Using both etches sequentially reduced the dielectric losses substantially compared to the untreated samples, by a factor of three, down to $\delta_{\mathrm{LP}}=5.2\cdot 10^{-7}$ or $Q_{\mathrm{LP}}=1.9$ M. Overall, our results demonstrate both the importance and the feasibility of targeted post-treatments on single chips, even within an industry-style fabrication environment.
Acknowledgements
The authors acknowledge helpful discussions with G. Huber, I. Tsitsilin, F. Haslbeck, C. Schneider, L. Koch and N. Bruckmoser from the Quantum Computing group at the Walther Meissner Institute. The authors also thank Z. Luo for his support with simulation. Additionally, we thank A. Reinholdt and his team from Fraunhofer ISC for the XPS measurements. We also appreciate M. Hahn and M. König for their help and valuable discussions in process development and the whole Fraunhofer EMFT clean room staff for the professional fabrication, especially Luca Rommeis.
This work was funded by the Munich Quantum Valley (MQV) – Consortium Scalable Hardware and Systems Engineering (SHARE), funded by the Bavarian State Government with funds from the Hightech Agenda Bavaria, the Munich Quantum Valley Quantum Computer Demonstrator - Superconducting Qubits (MUNIQC-SC) 13N16188, funded by the Federal Ministry of Education and Research, Germany, and the Open Superconducting Quantum Computers (OpenSuperQPlus) Project - European Quantum Technology Flagship.

### v2

@@
 Aluminum remains the central material for superconducting qubits, and considerable effort has been devoted to optimizing its deposition and patterning for quantum devices.
-However, while post-processing of Nb- and Ta-based resonators has been widely explored, primarily focusing on oxide removal using buffered oxide etch (BOE), post-treatment strategies for Al resonators remain underdeveloped.
+However, post-processing strategies focused on oxide removal of Nb- and Ta-based resonators using buffered oxide etch (BOE), which can not be used for Al.
 This challenge becomes particularly relevant for industry-scale fabrication with multi-chip bonding, where delays between sample preparation and cooldown require surface treatments that preserve low dielectric loss during extended exposure to ambient conditions.
@@
 Passivation using self-limiting oxygen and fluorine chemistries was evaluated utilizing different plasma processes.
-Remote oxygen plasma treatment reduced dielectric losses, in contrast to direct plasma, likely due to additional ashing of residual resist despite the formation of a thicker oxide layer on both Si and Al surfaces.
-A fluorine-based plasma process was developed that passivated the Al surface with fluorine for subsequent BOE treatment.
-However, increasing fluorine incorporation in the aluminum oxide correlated with higher loss, identifying fluorine as an unsuitable passivation material for Al resonators.
-Finally, selective oxide removal using HF vapor and phosphoric acid was assessed for surface preparation.
-HF vapor selectively etched SiO2 while preserving Al2O3, whereas phosphoric acid exhibited the opposite selectivity.
-Sequential application of both etches yielded dielectric losses as low as $\delta_{\mathrm{LP}}=5.2\times 10^{-7}$ ($Q\mathrm{i}\approx 1.9\,\mathrm{M}$) in the single photon regime, demonstrating a promising pathway for robust Al-based resonator fabrication.
+Remote oxygen plasma treatment reduced dielectric losses, in contrast to direct oxygen plasma.
+A fluorine-based plasma process was developed that passivated the Al surface for subsequent BOE treatment.
+However, the fluorine content in the surface resulted in higher loss, identifying fluorine as an unsuitable passivation material for Al resonators.
+Above all, selective oxide removal using HF vapor and phosphoric acid yielded dielectric losses as low as $\delta_{\mathrm{LP}}=5.2\times 10^{-7}$ ($Q\mathrm{i}\approx 1.9\,\mathrm{M}$) in the single photon regime.
+Selective oxide removal provides a promising pathway for robust Al-based qubit fabrication, as it preserves low dielectric losses for a 24-hour delay before cooldown.
 I Introduction
@@
 While Nb and Ta start to replace Al in the fabrication of resonators due to their favorable properties [33, 4], Al continues to be indispensable for Josephson junction electrodes, making it the most critical component of superconducting qubits.
-Significant efforts have therefore been devoted to optimize Al deposition and patterning for superconducting quantum circuits [3, 11].
-While post-processing strategies for Nb and Ta resonators, typically aimed to improve resonator quality with surface cleaning and oxide removal [2, 24, 10] using buffered oxide etch (BOE)[35, 1, 36], have been extensively studied, those techniques can not easily be applied to Al metallization.
+Substantial efforts have therefore been devoted to optimize Al deposition and patterning for superconducting quantum circuits [3, 11].
+Post-processing strategies for Nb and Ta resonators using buffered oxide etch (BOE)[35, 1, 36] typically aim to improve resonator quality with surface cleaning and oxide removal [2, 24, 10].
+However, those techniques can not easily be applied to Al metallization.
 Systematic investigations of post-treatments for Al resonators [23, 37, 6] have proposed different strategies to address existing limitations, which are particularly important because dielectric losses associated with native and process-induced oxides on aluminum or silicon can limit coherence in superconducting circuits [12, 27].
@@
 Since oxides are known to host two-level systems (TLS) [25, 27, 12], removing them is expected to reduce dielectric losses.
-Our study demonstrated the effectiveness of surface preparation for enhancing the performance of Al-based superconducting resonators.
+Our study demonstrated that a well designed surface preparation with plasma and etch processes enhance the performance of Al-based superconducting resonators for quantum applications.
 II Analytical Methods
@@
 The chip design consists of nine coplanar waveguides (CPW) resonators (gap: 6 µm, width: 10 µm) with evenly spaced frequencies between 4 GHz and 5.6 GHz, all capacitively coupled to a common feedline.
-The ground plane surrounding the resonators is patterned with vortex-pinning holes to mitigate quasiparticle tunneling [27].
+The ground plane surrounding the resonators is patterned with vortex-pinning holes [26].
 The measurement procedure is as follows: after locating the resonance peaks with a coarse transmission sweep, the complex scattering parameter $S_{21}$ is recorded.
@@
 However, the high-power (290 W) direct plasma process increased $\tilde{\delta}_{\mathrm{LP}}$ beyond the upper quartile of the untreated reference.
-The elevated ion energies likely damage both the Si and Al surfaces, thereby increasing the dielectric losses.
-(AFM pictures pending)
+The elevated ion energies likely damage both the Si and Al surfaces, thereby increasing the dielectric losses, making this process unsuited for our post processing purpose.
 IV-B Passivation via Fluorination
@@
 Finally, selective post-etching with HF and phosphoric acid was employed to remove SiO2 and Al2O3 from the resonator surfaces.
-Using both etches sequentially reduced the dielectric losses substantially compared to the untreated samples, by a factor of three, down to $\delta_{\mathrm{LP}}=5.2\cdot 10^{-7}$ or $Q_{\mathrm{LP}}=1.9$ M.
+Using both etches sequentially reduced the dielectric losses substantially compared to the untreated samples, by a factor of three, down to $\delta_{\mathrm{LP}}=5.2\cdot 10^{-7}$ or $Q_{\mathrm{LP}}=1.9$ M and is therefore very suitable for the post processing of superconducting Al-based qubits.
 Overall, our results demonstrate both the importance and the feasibility of targeted post-treatments on single chips, even within an industry-style fabrication environment.